OpenStudy (anonymous):

[Part 1] Show all work in simplifying 2x^2+8x-42/x^2+12x+35 (Be sure to list restrictions). [Part 2] Use complete sentences to explain your process.

OpenStudy (tkhunny):

#1) Completely factor numerator and denominator.

OpenStudy (anonymous):

please teach me how to do that @tkhunny

OpenStudy (anonymous):

2x^2+8x-42 -------------- x^2+12x+35

OpenStudy (tkhunny):

You should know how to factor. (2x^2 + 8x - 42) = 2(x^2 + 4x - 21) = 2(x + ___)(x - ___) Fill in the blanks.